40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9211     Entry time: Sun Oct 6 23:50:14 2013
Author: rana 
Type: Configuration 
Category: SUS 
Subject: MC filters adjusted 
  1. Found the low pass filters OFF in a couple of the MC SUS damping loops. This allows injection of OSEM noise all the way up to ~100 Hz. I deleted unused ones and turned on the 'Cheby' for all SUS DOFs: cheby1("LowPass",2,0.1,3)cheby1("LowPass",6,1,12)gain(1.13501)
  2. Tuned the Bounce/Roll filters to catch the bounce and roll modes for all 3 MC SUS (based on the Mech Res Wiki page). These are now engaged on ALL MC SUS DOFs. They have been deleted from the ASCPIT/YAW filter banks and moved into the WFS screen into the MC1,2,3 filter banks there to be more in line with our knowledge of multi loop instability from notches in individual loops:ellip("BandStop",4,1,40,15.9,17.2)ellip("BandStop",4,1,40,23.5,24.7)gain(1.25893)
ELOG V3.1.3-